How moles are present in 0.5 moles of a mercury in stp

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
@ chintu 
i think the ques u asked is wrong as its the no of molecule present and not the moles . 
i am assuming it to be molecule here , 
the molar mass of hg is 200 U 
which means 1 mole of hg = 200 g 
1 mole = 200 g 
0.5 mole = 200 *0.5 = 100 g 
hence , the mass of 0.5 mol of hg is 100g 
1 mole = 6.022 * 10 ^ 23 molecules of hg 
so, 0.5 moles = 0.5 * 6.022*10^23= 3.011*10 ^ 23 
HOPE IT CLEARS NOW ..